bobbiejames:

jackrabbits are cabable of reaching speeds up to 40 miles per hour. how fast is this in feet per second? (round to the nearest whole number

Mercury:

\[\frac{ 40~miles }{ ~hour }\frac{ 1~hour }{ 3600~seconds }\frac{ 5280~feet }{ 1~mile }\] dimensional analysis. notice how miles and hours cancel out and you're left with feet + seconds